OpenStudy (anonymous):

Simplify. (x3/5)^2/3

OpenStudy (mertsj):

What do you do with the exponents when raising to a power?

OpenStudy (anonymous):

The quantity of x to the three fifths power raised to the two thirds power

OpenStudy (mertsj):

\[(x ^{\frac{3}{5}})^{\frac{2}{3}}\]

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[(a^m)^n=a ^{m \times n}\]

OpenStudy (mertsj):

What do you do with the exponents when raising to a power?

OpenStudy (mertsj):

Are you there? Are you going to answer?

OpenStudy (anonymous):

Multiple choice a.) x9/125 b.) x5/8 c.) x2/5 d.) x6/15

OpenStudy (anonymous):

Simplify to a.) x9/125

OpenStudy (mertsj):

For the third and final time: What do you do with the exponents when raising to a power?

OpenStudy (anonymous):

Multiply

OpenStudy (mertsj):

What is 3/5 times 2/3?

OpenStudy (mertsj):